# Nontrivial twisted bilayer chiral excitonic systems: Förster coupling and related Hall effect

### v1

@@ II.1 The time-reversal even Hall effect (TREHE)
where two layers are divided into the system layer (top layer) and environment layer (bottom layer). $\boldsymbol{j}^{\mathrm{sys}}=-\boldsymbol{j}^{\mathrm{env}}$ preserves by the Onsager relation [43]. $f_{n}\left(\boldsymbol{k}\right)\approx f_{n}^{0}\left(\boldsymbol{k}\right)-\tau\frac{df_{n}^{0}\left(\boldsymbol{k}\right)}{dt}$ is the off-equilibrium distribution function expanded in the first order of relaxation time $\tau$, with $f_{n}^{0}\left(\boldsymbol{k}\right)=1/\left[\exp\left(\frac{\varepsilon_{n}\left(\boldsymbol{k}\right)-\mu}{k_{B}T}\right)-1\right]$ is the equilibrium Bose-Einstein distribution function. $\boldsymbol{v}_{n}^{\mathrm{sys/env}}\left(\boldsymbol{k}\right)=\left\langle u_{n}\left(\boldsymbol{k}\right)\right|\frac{1}{2}\left\{\boldsymbol{\hat{v}},\hat{P}^{\mathrm{sys/env}}\right\}\left|u_{n}\left(\boldsymbol{k}\right)\right\rangle$ is the velocity projection on system/environment layer of band $n$, where $\hat{P}^{\mathrm{sys/env}}=\left(1\pm\hat{\sigma}_{z}\right)/2$ is the projection operator onto the system/environment layer, with the Pauli matrix $\hat{\sigma}_{z}$ being the out-of-plane layer-pseudospin (LPS) operator operating in the layer index subspace [49]. $\boldsymbol{\hat{v}}=\frac{\partial H_{\mathrm{ex}}\left(\boldsymbol{k}\right)}{\hbar\partial\boldsymbol{k}}$ is the total velocity operator.
If we consider that there is an inhomogeneous distribution of the temperature $T$ in real space and a general mechanical in-plane force $\boldsymbol{F}$ as $\boldsymbol{F}=\hbar\boldsymbol{\dot{k}}$. After some tedious derivations [36], we can get the time-reversal even Hall conductivity of the system as $\sigma_{H}^{\mathrm{sys}}=\tau\mathcal{V}/\hbar$, while $\alpha_{H}^{\mathrm{sys}}=\tau\mathcal{V}_{\mathrm{Ner}}/\hbar$ represents the time-reversal even Nernst conductivity, where
@@
with $\omega_{n}\left(\boldsymbol{k}\right)=\frac{1}{2}\left[\nabla_{\boldsymbol{k}}\times\boldsymbol{v}_{n}^{\mathrm{sys}}\left(\boldsymbol{k}\right)\right]$ means the $\boldsymbol{k}$-space vorticity of the layer current [43].
In TMDs without strain, although the chiral symmetry in real space is preserved, the rotational symmetry of the monolayer TMDs ensures that the effective excitonic Hamiltonian is independent of the twisted angle $\theta$ [39]. This reinstates the mirror symmetry and leads to isotropic dispersions, as shown in Fig. 1(d). Consequently, the TREHE is forbidden. However, if strain is applied to the bottom layer of twisted bilayer TMDs, strain transfer between the two layers [47, 48] will result in a twisted-angle-dependent strain in the top layer [47, 48, 36]. These two inequivalent strains in the different layers can maintain chiral symmetry in the excitonic Hamiltonian, leading to a obvious TREHE, as shown in Fig. 2(b). Both $\sigma_{H}^{\mathrm{sys}}$ and $\alpha_{H}^{\mathrm{sys}}$ are periodic functions of $\theta$, with a period of $\pi/2$ due to $\theta$-dependent exchange coupling between excitons in different valleys of top layer, indicated by the product of two velocities in Fig. 3(a) [50].
In BP, due to the high anisotropy of the material, the chiral symmetry in real space is retained in the excitonic Hamiltonian, leading to anisotropic dispersions and a non-zero TREHE, as shown in Figs. 1(c) and 2(c). Here, $\sigma_{H}^{\mathrm{sys}}$ and $\alpha_{H}^{\mathrm{sys}}$ are also periodic functions of $\theta$, but with a period of $\pi$ due to the $C_{2}$ symmetry of the product of two velocities within the integral, as illustrated in Fig. 3(b). The chemical potential $\mu$ in Fig. 2 is chosen such that $\min\varepsilon_{1}-\mu\approx 0.005\,\mathrm{eV}$, where $\varepsilon_{1}$ is the ground energy of the excitonic Hamiltonian, and $\min\varepsilon_{1}$ in both materials is independent of $\theta$. Based on this consideration, the changes of $\sigma_{H}^{\mathrm{sys}}$ and $\alpha_{H}^{\mathrm{sys}}$ with respect to $T$ can be easily understand from the $T$ dependence of $f_{n=1}^{0}$ since it obviously increases as $T$ rises. This implies that the TREHE for excitons in chiral bilayer structures could be measurable at room temperature.
@@
Figure 4: (color online) (a) Schematic illustration of the crossed nonlinear dynamical Hall effect for excitons in twisted bilayer systems. (b) Numerical calculation of $\sigma_{H}$ and $\alpha_{H}$ (Eq. 30) for effective excitonic Hamiltonian $H_{\mathrm{TMD}}$ (Eq. 17) when strain is applied, with respect to the twisted angle $\theta$ and the temperature $T$, respectively. $\mu\approx 1.69\,\mathrm{eV}$. (c) The similar plot for effective excitonic Hamiltonian $H_{\mathrm{BP}}$ (Eq. 26), with $\mu\approx 1.98\,\mathrm{eV}$ with a gap $\varepsilon_{d}=0.02\,\mathrm{eV}$ induced by the physical control of layers such as strain. $F_{\perp}^{0}=0.1\,\mathrm{eV/nm}$, $\omega/2\pi=0.1\,\mathrm{THz}$ for all plots [44].
II.2 The crossed nonlinear dynamical Hall effect (CNDHE)
According to previous theoretical work [44, 45], this effect can be described from the similar starting point by calculating the current density of excitons in the intrinsic response
@@ II.2 The crossed nonlinear dynamical Hall effect (CNDHE)
is intrinsic to the band structure, measuring the $\boldsymbol{k}$-space curl of the interlayer Berry connection polarizability (BCP)
@@
over the occupied states [44, 45]. Here, $p_{z}=\Delta z\hat{\sigma}_{z}$ is the interlayer dipole moment. Although CNDHE remains forbidden in TMDs without strain since the high symmetry of the related excitonic Hamiltonian, it can emerge after applying strain, as demonstrated in Fig. 4(b), which shows a similar dependence to that observed in the TREHE calculation. This behavior is dominated by the product of the band velocity and BCP, as shown in Fig. 3(c) [50]. In BP, however, a significant value of CNDHE appears at specific twisted angle $\theta$ (Fig. 4(c)), corresponding to the distribution of BCP with respect to $\theta$, as shown in Fig. 3(d). The periodicity dependent on $\theta$ is $\pi$ since only $C_{2}$ symmetry are left in $H_{\mathrm{BP}}$. Compared to the TREHE, the CNDHE is more dependent on the material’s anisotropy and the specific twisted angle $\theta$, as evidenced by the numerical calculations in Fig. 3(d). This finding may suggest new opportunities for applications of BP in valley exciton optoelectronics.
@@ Acknowledgements.
C. L. would like to thank D. W. Zhai, B. Fu, and B. B for useful discussions. This work is supported by the National Key R&\&&D Program of China (2020YFA0309600), Research Grant Council of Hong Kong SAR (AoE/P-701/20, HKU SRFS2122-7S05), and New Cornerstone Science Foundation.

### v2

@@ II.1 The time-reversal even layer Hall (Nernst) counter flow (TREHCF)
where two layers are divided into the system layer (top layer) and environment layer (bottom layer). $\boldsymbol{j}^{\mathrm{sys}}=-\boldsymbol{j}^{\mathrm{env}}$ preserves by the Onsager relation [46]. $f_{n}\left(\boldsymbol{k}\right)\approx f_{n}^{0}\left(\boldsymbol{k}\right)-\tau\frac{df_{n}^{0}\left(\boldsymbol{k}\right)}{dt}$ is the off-equilibrium distribution function expanded in the first order of relaxation time $\tau$, with $f_{n}^{0}\left(\boldsymbol{k}\right)=1/\left[\exp\left(\frac{\varepsilon_{n}\left(\boldsymbol{k}\right)-\mu}{k_{B}T}\right)-1\right]$ is the equilibrium Bose-Einstein distribution function. $\boldsymbol{v}_{n}^{\mathrm{sys/env}}\left(\boldsymbol{k}\right)=\left\langle u_{n}\left(\boldsymbol{k}\right)\right|\frac{1}{2}\left\{\boldsymbol{\hat{v}},\hat{P}^{\mathrm{sys/env}}\right\}\left|u_{n}\left(\boldsymbol{k}\right)\right\rangle$ is the velocity projection on system/environment layer of band $n$, where $\hat{P}^{\mathrm{sys/env}}=\left(1\pm\hat{\sigma}_{z}\right)/2$ is the projection operator onto the system/environment layer, with the Pauli matrix $\hat{\sigma}_{z}$ being the out-of-plane layer-pseudospin operator operating in the layer index subspace [52]. $\boldsymbol{\hat{v}}=\frac{\partial H_{\mathrm{ex}}\left(\boldsymbol{k}\right)}{\hbar\partial\boldsymbol{k}}$ is the total velocity operator.
If we consider that there is an inhomogeneous distribution of the temperature $T$ in real space and a general mechanical in-plane force $\boldsymbol{F}$ as $\boldsymbol{F}=\hbar\boldsymbol{\dot{k}}$. After some tedious derivations [36], we can get the time-reversal even Hall conductivity for the layer counter flow of the system as $\sigma_{H}^{\mathrm{sys}}=\tau\mathcal{V}/\hbar$, while $\alpha_{H}^{\mathrm{sys}}=\tau\mathcal{V}_{\mathrm{Ner}}/\hbar$ represents the time-reversal even Nernst conductivity for the layer counter flow, where
@@
with $\omega_{n}\left(\boldsymbol{k}\right)=\frac{1}{2}\left[\nabla_{\boldsymbol{k}}\times\boldsymbol{v}_{n}^{\mathrm{sys}}\left(\boldsymbol{k}\right)\right]$ means the $\boldsymbol{k}$-space vorticity of the layer current [46], exhibiting the nontrivial quantum geometric properties characteristic of the layer hybridized wavefunctions under the chiral symmetry.
In TMDs without strain, although the chiral symmetry in real space is preserved, the rotational symmetry of the monolayer TMDs ensures that the effective exciton Hamiltonian is independent of the twisted angle $\theta$ [41]. This reinstates the mirror symmetry and leads to isotropic dispersions, as shown in Fig. 1(c). Consequently, the TREHCF is forbidden. However, if strain is applied to the bottom layer of twisted bilayer TMDs, strain transfer between the two layers [50, 51] will result in a twisted-angle-dependent strain in the top layer [50, 51, 36]. These two inequivalent strains in the different layers can break the rotational symmetry in the exciton Hamiltonian, leading to the excitonic chirality and a obvious TREHCF, as shown in Fig. 2(b). Both $\sigma_{H}^{\mathrm{sys}}$ and $\alpha_{H}^{\mathrm{sys}}$ are periodic functions of $\theta$, with a period of $\pi/2$ dominated by the quantum geometric properties of $\omega_{n}\left(\boldsymbol{k}\right)$, as shown from the product of two velocities in Fig. 3(a) [53].
In BP, due to the high anisotropy, the excitonic chirality is naturally satisfied, leading to anisotropic exciton dispersions and a non-zero TREHCF, as shown in Figs. 1(d) and 2(c). Here, $\sigma_{H}^{\mathrm{sys}}$ and $\alpha_{H}^{\mathrm{sys}}$ are also periodic functions of $\theta$, but with a period of $\pi$ due to the $C_{2}$ symmetry of the product of two velocities within the integral, as illustrated in Fig. 3(b). The chemical potential $\mu$ in Fig. 2 is chosen such that $\min\varepsilon_{1}-\mu\approx 0.005\,\mathrm{eV}$, where $\varepsilon_{1}$ is the ground energy of the exciton Hamiltonian, and $\min\varepsilon_{1}$ in both materials is independent of $\theta$. Based on this consideration, the changes of $\sigma_{H}^{\mathrm{sys}}$ and $\alpha_{H}^{\mathrm{sys}}$ with respect to $T$ can be easily understand from the $T$ dependence of $f_{n=1}^{0}$ since it obviously increases as $T$ rises. This implies that the TREHCF for excitons in chiral bilayer structures could be measurable at room temperature.
@@
Figure 4: (color online) (a) Schematic illustration of the crossed nonlinear dynamical Hall (Nernst) effect for excitons in twisted bilayer systems. (b) Numerical calculation of $\sigma_{H}$ and $\alpha_{H}$ (Eq. 30) for effective exciton Hamiltonian $H_{\mathrm{TMD}}$ (Eq. 17) when strain is applied, with respect to the twisted angle $\theta$ and the temperature $T$, respectively. $\mu\approx 1.69\,\mathrm{eV}$. (c) The similar plot for effective exciton Hamiltonian $H_{\mathrm{BP}}$ (Eq. 26), with $\mu\approx 1.98\,\mathrm{eV}$ with a gap $\varepsilon_{d}=0.02\,\mathrm{eV}$ induced by the physical control of layers such as strain. $F_{\perp}^{0}=0.1\,\mathrm{eV/nm}$, $\omega/2\pi=0.1\,\mathrm{THz}$ for all plots [47].
II.2 The crossed nonlinear dynamical Hall (Nernst) effect (CNDHE)
According to previous theoretical works [47, 48], this effect can be described from the similar starting point by calculating the current density of excitons in the intrinsic response
@@ II.2 The crossed nonlinear dynamical Hall (Nernst) effect (CNDHE)
is intrinsic to the band structure. Here, the unusual quantum geometric properties induced by the chiral symmetric layer hybridized wavefunctions is measured by the $\boldsymbol{k}$-space curl of the interlayer Berry connection polarizability (BCP)
@@
over the occupied states [47, 48]. Here, $p_{z}=\Delta z\hat{\sigma}_{z}$ is the interlayer dipole moment. Although CNDHE remains forbidden in TMDs without strain since the rotational symmetry of the related exciton Hamiltonian, it can emerge after applying strain, as demonstrated in Fig. 4(b), which shows a similar dependence to that observed in the TREHCF calculation. This behavior is dominated by the quantum geometric properties of $\left(\nabla_{\boldsymbol{k}}\times\boldsymbol{\mathcal{G}}^{n}\left(\boldsymbol{k}\right)\right)_{z}$, as shown in Fig. 3(c) [53]. In BP, however, a significant value of CNDHE appears at specific twisted angle $\theta$ (Fig. 4(c)), corresponding to the distribution of BCP with respect to $\theta$, as shown in Fig. 3(d). The periodicity dependent on $\theta$ is $\pi$ since only $C_{2}$ symmetry are left in $H_{\mathrm{BP}}$. Compared to the TREHCF, the CNDHE is more dependent on the material’s anisotropy and the specific twisted angle $\theta$, as evidenced by the numerical calculations in Fig. 3(d). This finding may suggest new opportunities for applications of BP in exciton-based optoelectronics.
@@ Acknowledgements.
C. L. would like to thank D. W. Zhai, B. Fu, C. Xiao, and B. B for useful discussions. This work is supported by the National Key R&\&&D Program of China (2020YFA0309600), Research Grant Council of Hong Kong SAR (AoE/P-701/20, HKU SRFS2122-7S05), and New Cornerstone Science Foundation.